OpenStudy (anonymous):

Use Logarithms to solve the equation 7^x = 2^(x+1) giving the value of x to 3sf currently doing core 2 mathematics and am stuck on this one. any help would be appreciated

OpenStudy (anonymous):

\[7^x=2^{x+1}\]

OpenStudy (anonymous):

alright, where are you stuck at? The general application of logarithmns? Or does it confuse you because you have unlike bases?

OpenStudy (anonymous):

7^x = 2^(x+1) I take logs of both sides ending up with xln7=(x+1)ln2 where to go from here?

OpenStudy (anonymous):

very good so far, distribute the right hand side, regular distribution property and then save the linear equation for x.

OpenStudy (anonymous):

so is that xln7 = xln2 + ln2

OpenStudy (anonymous):

exactly, now try to solve for x, getting it on the same side and factor it out (distribution property in reverse -> factorisation)

OpenStudy (anonymous):

xln7 - xln 2 = ln2 x(ln7-ln2) = ln2 x(ln3.5) = ln2 x=ln2 over ln3.5

OpenStudy (anonymous):

there you go.

OpenStudy (anonymous):

xln7 = xln2 + ln2 is what i was struggling with. thankyou for your help

OpenStudy (anonymous):

you're very welcome!